# Relative Gieseker’s problem on $F$-divided bundles

### v1

@@ 3.2 The general case
In the proofs below we use the following important generalization of [EM10, Proposition 2.3] that is contained in [LZ25, Proposition 6.3].
@@
Let $f:X\to Y$ be a surjective morphism of normal projective $k$-varieties. If $f_{*}:\pi^{\sf{\acute{e}t}}_{1}(X)\to\pi^{\sf{\acute{e}t}}_{1}(Y)$ is surjective then $f_{*}:\pi^{F\sf{\textendash div}}_{1}(X)\to\pi^{F\sf{\textendash div}}_{1}(Y)$ is faithfully flat.
@@
is an isomorphism. By Proposition 3.7 we can assume that for all $j$ both $E_{1,j}$ and $E_{2,j}$ are slope (or even Gieseker) semistable. So the required assertion follows from Corollary 3.4.
By [DM82, Proposition 2.21] it is sufficient to show that for any ${\mathbb{E}}\in{\sf Vect}^{\sf perf}(Y)$ and any subobject ${\mathbb{G}}\subseteq f^{*}{\mathbb{E}}$, $\mathbb{G}$} is contained in the essential image of $f^{*}$. By Proposition 3.7 we can assume that $\mathbb{E}$} admits a filtration by subobjects so that the quotients ${\mathbb{F}}_{1},...,{\mathbb{F}}_{m}$ are $F$-divided bundles on $Y$ with the property that all $F_{j,n}$ are slope stable with numerically vanishing Chern classes. Then by Lemma 3.5 $f^{*}{\mathbb{E}}$ is a successive extension of the $F$-divided bundles $f^{*}{\mathbb{F}}_{j}\in{\sf Vect}^{{\sf perf}}(X)$ with all $f^{*}F_{j,n}$ slope stable with numerically vanishing Chern classes. In particular, all $f^{*}{\mathbb{F}}_{j}$ are simple objects of ${\sf Vect}^{{\sf perf}}(X)$. Since the category ${\sf Vect}^{{\sf perf}}(X)$ is Tannakian, an easy induction on the rank of $\mathbb{G}$} shows that $\mathbb{G}$} is an extension of some of the $F$-divided bundles $f^{*}{\mathbb{F}}_{1},...,f^{*}{\mathbb{F}}_{m}$ and hence it is contained in the essential image of $f^{*}$. ∎
@@
In case $f$ is a finite genuinely ramified morphism between smooth projective $k$-varieties the above theorem was claimed in [BKP25, Remark 5.1], who checked that in this case $f^{*}:{\sf Vect}^{{\sf perf}}(Y)\to{\sf Vect}^{{\sf perf}}(X)$ is fully faithful. It is not clear to the author how they planned to prove the second condition needed for application of [DM82, Proposition 2.21].
@@
If $f_{*}:\pi^{\sf{\acute{e}t}}_{1}(X)\to\pi^{\sf{\acute{e}t}}_{1}(Y)$ is surjective then $f^{*}:{\sf Vect}^{{\sf perf}}(Y)\to{\sf Vect}^{{\sf perf}}(X)$ is a fully faithful exact functor. However, the essential image of $f^{*}$ does not need to be a Serre subcategory. For example, if $X$ is an ordinary elliptic curve and $f$ is the map of $X$ to a point then the essential image of $f^{*}$ consists of trivial objects but an extension of ${\mathbbm{1}}_{X}$ by ${\mathbbm{1}}_{X}$ need not be trivial as ${\mathop{{\rm Ext\,}}}^{1}({\mathbbm{1}}_{X},{\mathbbm{1}}_{X})=H^{1}(X,{\mathcal{O}}_{X})\neq 0$ (see, e.g., the proof of [dS07, Theorem 15]).
@@
Then we construct $S$-flat models $A_{S}\subset R(r,X_{S},x_{S})$ for $A$ and $A_{S}^{\prime}\subset R(r,X^{\prime}_{S},x^{\prime}_{S})$ for $A^{\prime}$. We have a dominant rational map of $S$-schemes $\varphi_{S}:A^{\prime}_{S}\dashrightarrow A_{S}$ extending $\varphi$ and defined by pullback via the relative Frobenius morphism $F_{X_{S}/S}:X_{S}\to X^{\prime}_{S}$. Shrinking $S$ if necessary we can assume that the restriction $\varphi_{s}:A^{\prime}_{s}\dashrightarrow A_{s}$ is a dominant rational map for all closed points $s$ of $S$. For such $s$ we let $\varphi_{s}^{(i)}$ denote the $i$th Frobenius twist of $\varphi_{s}$ and set $m_{s}=(\kappa(s):{\mathbb{F}}_{p})$. Since $\varphi_{s}$ is defined by $(E,\beta)]\to[F_{X_{s}/s}^{*}E,F_{\kappa(s)}^{*}\beta]$, the composition $\varphi_{s}^{(m_{s}-1)}\circ...\circ\varphi_{s}^{(1)}\circ\varphi_{s}$ defines a rational endomorphism $A_{s}\dashrightarrow A_{s}$ for any closed point $s\in S$. Now (as in [EM10]) Hrushovski’s theorem [Hru04, Corollary 1.2] implies that set of closed points of $A_{s}$, which are periodic for this rational endomorphism is dense in $A_{s}$. Such points correspond to framed pairs $(G,\beta_{G})]$, where $G$ is a geometrically Gieseker stable vector bundle on $X_{s}$ such that for some $m\geq 1$ we have an isomorphism $(F^{m}_{X_{s}})^{*}G\simeq G$. If $\bar{s}$ is a geometric point lying over $s$ then $G_{\bar{s}}$ on $X_{\bar{s}}$ gives by [Kat73, Proposition 4.1.1] a representation of $\pi^{\sf{\acute{e}t}}_{1}(X_{\bar{s}})$.
Let us consider $X\times_{Y}X$ with two projections $p_{1}$ and $p_{2}$ onto $X$. Similarly, let us define $p_{1\bar{s}}$ and $p_{2\bar{s}}$. By [SGA03, Exposé IX, Corollaire 5.6] and [SGA03, Exposé IX, Théorème 4.12] and our assumption, it follows that the two homomorphisms $p_{1*},p_{2*}:\pi^{\sf{\acute{e}t}}_{1}(X\times_{Y}X)\to\pi^{\sf{\acute{e}t}}_{1}(X)$ coincide. By [Aut, Lemma 0C0K] we have a commutative diagram
@@
By [Aut, Lemma 0C0P], the vertical specialization maps are surjective, so we have $p_{1\bar{s}*}=p_{2\bar{s}*}$. But then [SGA03, Exposé IX, Corollaire 5.6] and [SGA03, Exposé IX, Théorème 4.12] imply that the map $f_{*}:\pi^{\sf{\acute{e}t}}_{1}(X_{\bar{s}})\to\pi^{\sf{\acute{e}t}}_{1}(Y_{\bar{s}})$ is an isomorphism.
In this way, $G_{\bar{s}}$ induces a representation of $\pi^{\sf{\acute{e}t}}_{1}(Y_{\bar{s}})$, which by [Kat73, Proposition 4.1.1] gives rise to a vector bundle $G^{\prime}$ on $Y_{\bar{s}}$ such that $G_{\bar{s}}\simeq f_{\bar{s}}^{*}G^{\prime}$. Since the framing $\beta_{G_{\bar{s}}}:x_{\bar{s}}^{*}G_{\bar{s}}\simeq{\mathcal{O}}_{\kappa(\bar{s})}^{\oplus r}$ descends to a framing of $G^{\prime}$ at $f_{\bar{s}}(x_{\bar{s}})$, the rational map $f_{s}^{*}:(f_{s}^{*})^{-1}(A_{s})\dashrightarrow A_{s}$, given by the pullback $(E,\beta)]\to[(f^{*}_{s}E,f_{s}^{*}\beta)]$, is dominant.
@@ Acknowledgements
The author was partially supported by Polish National Centre (NCN) contract number 2021/41/B/ST1/03741. The author would like to thank Lei Zhang for useful conversations and previous colaboration on the related project [LZ25].

### v2

@@ 3.2 The general case
In the proofs below we use the following important generalization of [9, Proposition 2.3] that is contained in [22, Proposition 6.3].
@@
Let $f:X\to Y$ be a surjective morphism of normal projective $k$-varieties. If $Y$ is smooth and $f_{*}:\pi^{\sf{\acute{e}t}}_{1}(X)\to\pi^{\sf{\acute{e}t}}_{1}(Y)$ is surjective then $f_{*}:\pi^{F\sf{\textendash div}}_{1}(X)\to\pi^{F\sf{\textendash div}}_{1}(Y)$ is faithfully flat.
@@
is an isomorphism. Indeed, by Proposition 3.7 we can assume that for all $j$ both $E_{1,j}$ and $E_{2,j}$ are slope (or even Gieseker) semistable. So the required assertion follows from the fact that for any slope semistable vector bundles $E_{1}$ and $E_{2}$ the canonical map
${\mathop{{\rm Hom}}}_{Y}(E_{1},E_{2})\to{\mathop{{\rm Hom}}}_{X}(f^{*}E_{1},f^{*}E_{2})$
is an isomorphism. This last isomorphism follows from Corollary 3.4 as $\pi^{\sf{\acute{e}t}}_{1}(X_{\sf reg})\to\pi^{\sf{\acute{e}t}}_{1}(X)$ is surjective, and hence $f$ is regularly genuinely ramified.
By [6, Proposition 2.21] to finish the proof it is sufficient to show that for any ${\mathbb{E}}\in{\sf Vect}^{\sf perf}(Y)$ and any subobject ${\mathbb{G}}\subseteq f^{*}{\mathbb{E}}$, $\mathbb{G}$} is contained in the essential image of $f^{*}$. By Proposition 3.7 we can assume that $\mathbb{E}$} admits a filtration by subobjects so that the quotients ${\mathbb{F}}_{1},...,{\mathbb{F}}_{m}$ are $F$-divided bundles on $Y$ with the property that all $F_{j,n}$ are slope stable with numerically vanishing Chern classes. Then by Lemma 3.6 $f^{*}{\mathbb{E}}$ is a successive extension of the $F$-divided bundles $f^{*}{\mathbb{F}}_{j}\in{\sf Vect}^{{\sf perf}}(X)$ with all $f^{*}F_{j,n}$ slope stable with numerically vanishing Chern classes. In particular, all $f^{*}{\mathbb{F}}_{j}$ are simple objects of ${\sf Vect}^{{\sf perf}}(X)$. Since the category ${\sf Vect}^{{\sf perf}}(X)$ is Tannakian, an easy induction on the rank of $\mathbb{G}$} shows that $\mathbb{G}$} is an extension of some of the $F$-divided bundles $f^{*}{\mathbb{F}}_{1},...,f^{*}{\mathbb{F}}_{m}$ and hence it is contained in the essential image of $f^{*}$. ∎
@@
In case $f$ is a finite genuinely ramified morphism between smooth projective $k$-varieties the above theorem was claimed in [3, Remark 5.1], who checked that in this case $f^{*}:{\sf Vect}^{{\sf perf}}(Y)\to{\sf Vect}^{{\sf perf}}(X)$ is fully faithful. It is not clear to the author how they planned to prove the second condition needed for application of [6, Proposition 2.21].
@@
If $f_{*}:\pi^{\sf{\acute{e}t}}_{1}(X)\to\pi^{\sf{\acute{e}t}}_{1}(Y)$ is surjective then $f^{*}:{\sf Vect}^{{\sf perf}}(Y)\to{\sf Vect}^{{\sf perf}}(X)$ is a fully faithful exact functor. However, the essential image of $f^{*}$ does not need to be a Serre subcategory. For example, if $X$ is an ordinary elliptic curve and $f$ is the map of $X$ to a point then the essential image of $f^{*}$ consists of trivial objects but an extension of ${\mathbbm{1}}_{X}$ by ${\mathbbm{1}}_{X}$ need not be trivial as ${\mathop{{\rm Ext\,}}}^{1}({\mathbbm{1}}_{X},{\mathbbm{1}}_{X})=H^{1}(X,{\mathcal{O}}_{X})\neq 0$ (see, e.g., the proof of [7, Theorem 15]).
@@
Then we construct $S$-flat models $A_{S}\subset R(r,X_{S},x_{S})$ for $A$ and $A_{S}^{\prime}\subset R(r,X^{\prime}_{S},x^{\prime}_{S})$ for $A^{\prime}$. We have a dominant rational map of $S$-schemes $\varphi_{S}:A^{\prime}_{S}\dashrightarrow A_{S}$ extending $\varphi$ and defined by pullback via the relative Frobenius morphism $F_{X_{S}/S}:X_{S}\to X^{\prime}_{S}$. Shrinking $S$ if necessary we can assume that the restriction $\varphi_{s}:A^{\prime}_{s}\dashrightarrow A_{s}$ is a dominant rational map for all closed points $s$ of $S$. For such $s$ we let $\varphi_{s}^{(i)}$ denote the $i$th Frobenius twist of $\varphi_{s}$ and set $m_{s}=(\kappa(s):{\mathbb{F}}_{p})$. Since $\varphi_{s}$ is defined by $(E,\beta)]\to[F_{X_{s}/s}^{*}E,F_{\kappa(s)}^{*}\beta]$, the composition $\varphi_{s}^{(m_{s}-1)}\circ...\circ\varphi_{s}^{(1)}\circ\varphi_{s}$ defines a rational endomorphism $A_{s}\dashrightarrow A_{s}$ for any closed point $s\in S$. Now (as in [9]) Hrushovski’s theorem [18, Corollary 1.2] implies that set of closed points of $A_{s}$, which are periodic for this rational endomorphism is dense in $A_{s}$. Such points correspond to framed pairs $(G,\beta_{G})]$, where $G$ is a geometrically Gieseker stable vector bundle on $X_{s}$ such that for some $m\geq 1$ we have an isomorphism $(F^{m}_{X_{s}})^{*}G\simeq G$. If $\bar{s}$ is a geometric point lying over $s$ then $G_{\bar{s}}$ on $X_{\bar{s}}$ gives by [20, Proposition 4.1.1] a representation of $\pi^{\sf{\acute{e}t}}_{1}(X_{\bar{s}})$.
Let us consider $X\times_{Y}X$ with two projections $p_{1}$ and $p_{2}$ onto $X$. Similarly, let us define $p_{1\bar{s}}$ and $p_{2\bar{s}}$. By [26, Exposé IX, Corollaire 5.6] and [26, Exposé IX, Théorème 4.12] and our assumption, it follows that the two homomorphisms $p_{1*},p_{2*}:\pi^{\sf{\acute{e}t}}_{1}(X\times_{Y}X)\to\pi^{\sf{\acute{e}t}}_{1}(X)$ coincide. By [1, Lemma 0C0K] we have a commutative diagram
@@
By Lemma 1.7, shrinking $S$ if necessary, we can assume that the vertical specialization maps are surjective and hence we have $p_{1\bar{s}*}=p_{2\bar{s}*}$. But then [26, Exposé IX, Corollaire 5.6] and [26, Exposé IX, Théorème 4.12] imply that the map $f_{*}:\pi^{\sf{\acute{e}t}}_{1}(X_{\bar{s}})\to\pi^{\sf{\acute{e}t}}_{1}(Y_{\bar{s}})$ is an isomorphism.
In this way, $G_{\bar{s}}$ induces a representation of $\pi^{\sf{\acute{e}t}}_{1}(Y_{\bar{s}})$, which by [20, Proposition 4.1.1] gives rise to a vector bundle $G^{\prime}$ on $Y_{\bar{s}}$ such that $G_{\bar{s}}\simeq f_{\bar{s}}^{*}G^{\prime}$. Since the framing $\beta_{G_{\bar{s}}}:x_{\bar{s}}^{*}G_{\bar{s}}\simeq{\mathcal{O}}_{\kappa(\bar{s})}^{\oplus r}$ descends to a framing of $G^{\prime}$ at $f_{\bar{s}}(x_{\bar{s}})$, the rational map $f_{s}^{*}:(f_{s}^{*})^{-1}(A_{s})\dashrightarrow A_{s}$, given by the pullback $(E,\beta)]\to[(f^{*}_{s}E,f_{s}^{*}\beta)]$, is dominant.
@@ Acknowledgements
The author was partially supported by Polish National Centre (NCN) contract number 2021/41/B/ST1/03741. The author would like to thank Lei Zhang for useful conversations and previous colaboration on the related project [22]. He would also like to thank Dario Weißmann for pointing out some errors in the first version of the paper.